Diffusion Morning Report 04/20/11

Ted Berg tberg at stanford.edu
Wed Apr 20 06:11:56 PDT 2011


Good morning All,
     Looking good this morning. B2H6 and Germane cylinders for Epi 
changed yesterday- will be on today. Noise filter installed on TEL.Feild 
service in yesterday to tweek 4160- need inputs from users. Tylan LTO 
had O2 flow issues. Thermco 2 had no TLC flow. No other issues at this time.

					
	Date: 	4/20/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*up* 	** 	*5.25/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*1.94/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*.62/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*11.23/30* 	** 	**
Tylan 11-LTO 	*yellow* 	*O2 flow issues* 	*16.9/30* 	** 	*2*
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	*
* 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi 	*yellow* 	*no PH3 -B2H6 and GeH4 cylinders changed* 	** 	** 	*1*
TEL Spa 	*up* 	*noise filter installed* 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor dead* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*A on line 446 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110420/6dd8abee/attachment.html>


More information about the diffusion mailing list